# LiFT: A Surprisingly Simple Lightweight Feature Transform for Dense ViT Descriptors

### v1

@@ 6.1 LiFT and Scale Invariance of Features
In this experiment, we demonstrate that LiFT intrinsically learns to generate features that are more scale-invariant. We use the Centered Kernel Alignment (CKA) metric [13, 27, 41], which can measure the similarity of a pair of feature maps even when they are different sizes. Using images in the SPair-71k training set, we re-scale each image to a range of different sizes and then extract features with DINO or DINO+LiFT and measure the CKA similarity between all input size pairings. As a baseline, we also compare with bilinearly upsampled DINO features. In Figure 4, we take four source scales and plot the CKA similarity with the features at all other scales. We see that LiFT greatly improves the inter-scale feature similarity for small input scales, and moderately improves the similarity for medium and larger scales. This shows that LiFT produces representations that are more scale-invariant than the base DINO features. Bilinear upsampling does provide a small improvement in CKA similarity across scales, though the benefit is far smaller than that of LiFT. This scale invariance property of LiFT is learned automatically, and likely comes from its multi-scale reconstruction objective. LiFT must learn to counteract the effect of input scale in order to map features from low resolution inputs to those of high resolution inputs. This property is desirable for dense tasks where objects appear at different scales.
6.2 Enhanced Self-Similarity Maps with LiFT
@@
To further improve our understanding of the dense features generated by LiFT, we visualize the self-similarity of the features in Figure 5. Base DINO S/16 yields $14\times 14$ features for a $224\times 224$ image. In our comparisons, we visualize these features alongside: a bilinearly interpolated upsampling of these same features to $28\times 28$ (DINO+Bilinear), DINO features for a $448\times 448$ input image, and finally DINO+LiFT features generated for a $224\times 224$ input image. The last three configurations all yield $28\times 28$ feature grids. To visualize these features, we select the center-most token feature for each of the maps and compute its similarity with all other features and visualize the similarity scores.
We show six feature similarity maps in Figure 5, and additional sample visualizations can be found in Appendix 0.E. We find that, qualitatively, the LiFT output gives a cleaner boundary and better highlights the content corresponding to the central patch. Having a high similarity to relevant regions and a clear boundary is beneficial for multiple localized downstream tasks. It also indicates that the DINO+LiFT features have better spatial awareness of object boundaries. Note that for row 2, where the central pixel corresponds to the background region, the similarity map highlights the background in the image, though it still appears that DINO+LiFT produces better similarity maps with sharper edges. In row 3, when there are multiple objects of the same type, DINO+LiFT better highlights the separate object instances. These results are qualitative, but they suggest that the LiFT-enhanced features have better content and boundary information than the base DINO features, which likely contributes to their improved performance in correspondence and segmentation tasks.
6.3 Variations in Backbone
One of the requirements of a feature densifying approach is that it should be easy to train on any backbone. To this end, we show that our approach consistently gives a performance gain with multiple different backbones: DINO (Table 8 row 1 vs. row 2), MoCo (Table 8 row 4 vs. row 5) and a fully-supervised ViT (Table 8 row 7 vs. row 8). This consistent improvement shows that LiFT can be trained in the exact same manner on differently trained ViTs without need for careful hyperparameter tuning. To verify that LiFT does not simply learn a bilinear interpolation, we apply LiFT modules trained on one backbone to the output of a different backbone. We show these results in Table 7 using the PCK@0.1 metric and $448\times 448$ images. It can be seen that when LiFT is applied to a different model than what was used to train it, the performance drops and is lower than not applying LiFT. This shows that LiFT learns a model-specific feature-densifying transform and not a simple interpolation.
@@ 6.3 Variations in Backbone
6.4 Repeated Application of the LiFT Module
In our base approach, the LiFT module is applied once to the ViT features to double their resolution. In this section, we also test the potential benefits of applying the LiFT block multiple times. To check this, we take the LiFT module and super-resolute the features twice by passing the output of first super-resolution again through the same LiFT network. We denote this approach as ‘2$\times$×LiFT’ in Table 8. We can see that in most cases the performance increases after recursively applying LiFT. This is especially true for the $56\times 56$, $112\times 112$, and $224\times 224$ resolutions. For $448\times 448$, there are a few places where it negligibly drops performance, but it still shows improvement for most cases.
@@ 7 Conclusion and Discussion
We have presented LiFT, a simple yet effective self-supervised Lightweight Feature Transform to boost the density of features of pretrained ViT backbones. This approach allows us to extract higher resolution spatial features from ViTs which can then be used for multiple dense downstream tasks. LiFT is task-agnostic and gives significant boosts in SPair Keypoint Correspondence, DAVIS Video Object Segmentation, Unsupervised Object Discovery, and COCO Detection and Segmentation. This benefit comes for a fraction of the computational cost compared with other densification methods. LiFT is a lightweight module that is easily trained with a self-supervised objective, and it is far cheaper than full backbone fine-tuning, as is done by other prior works. Through extensive experiments, we have shown that LiFT can be trained easily on any backbone and consistently leads to improved performance by generating better quality, higher-density features. We also show that LiFT can be applied to its own output in a recursive manner enabling good performance with even lower image resolutions. Finally, we show that our surprisingly simple method has several desirable emergent properties including scale-invariant features, and better object boundary maps. This makes LiFT a useful multipurpose tool for many potential downstream applications.
Acknowledgements. We would like to thank our colleagues Matthew Gwilliam and Pravin Nagar for their feedback on this work.
@@ 0.A.4 Ablation of Training Epochs
Appendix 0.B Additional Details for ViTDet+LiFT
For our experiments combining LiFT with ViTDet [28], we increase the size of our LiFT module to address the additional complexity of the task and backbone. To be consistent with ViTDet, we use an MAE-trained ViT-Base backbone instead of the ViT-Small used in our other primary experiments. Note that a standard ViT-Small model outputs feature maps with $384$ channels, while ViT-Base outputs $768$ channels. To handle the increased number of channels, we commensurately increase the number of channels in the layers of our LiFT module. We also add an additional convolutional block to the encoder segment. This larger LiFT module has a total of $7M$ parameters, as compared with the $1.2M$ parameter version used for smaller architectures. The ViTDet model used has $111M$ parameters, so our combined ViTDet+LiFT architecture has $118M$ parameters total. This is a $6.3\%$ increase in total parameters, which is similar to the relative percentage increase of the smaller LiFT version used with DINO S/16. Also, here we train LiFT on the COCO dataset in place of ImageNet. Because the COCO dataset is much smaller than ImageNet, we train on it for 100 epochs.
@@ Appendix 0.C Additional Backbones with LiFT
0.C.1 Performance Improvements with LiFT for Other Backbones
We further demonstrate the general utility of LiFT by applying it to several additional backbones, including Leopart [55] and several other DINO [5] ViTs, namely ViT-S/8, ViT-B/16, and ViT-B/8. The results are summarized in Table 11. LiFT shows consistent improvement for the various architectures and models across patch sizes (8 and 16), trainings (Leopart and DINO) and backbone sizes (Base and Small).
0.C.2 Cost vs. Performance Trade-off for Other Backbones
We extend the Performance vs. Compute Cost analysis from Section 5.2 to include both the MOCO and fully-supervised ViT backbones, as shown in Figure 6. We find that LiFT consistently boosts the performance of all three backbones at all FLOP allowances.
@@ 0.C.1 Performance Improvements with LiFT for Other Backbones
Appendix 0.D Additional Metrics
We present results for additional metrics on the SPair-71k [32] and DAVIS [33] datasets. For SPair-71k, we additionally report the values for PCK@0.01 in Table 12. For DAVIS, we also report the J Mean and F Mean in Table 13. We present these results alongside the previously reported metrics for completeness. For SPair-71k, the PCK@0.01 metric demands the most precision, and the additional feature resolution provided by LiFT provides consistent improvements. For DAVIS, both the J Mean and F Mean are also consistently improved by adding LiFT.
Appendix 0.E Additional Similarity Map Samples
In Section 6.2, we found that the feature self-similarity maps for DINO+LiFT more clearly and sharply outline the central object in an image. To further highlight this, we provide a zoomed-in comparison of the difference between DINO+LiFT and DINO+Bilinear upscaling in Figure 7. We can see that DINO with Bilinear upsampling highlights the main object, but the outline is hazier and less precise due to the smoothing of the features. Meanwhile, the upscaled feature map produced by LiFT better respects object contours and produces a much sharper feature self-similarity map.
Finally, we provide additional samples further showing the benefits of LiFT for self-similarity maps, as shown in Figure 8. In rows 1 to 8 (left), we show samples with single central objects of differing shapes and sizes. We see that the feature self-similarity maps for DINO+LiFT more uniformly fill the foreground object region, and have less noisy correlations with background regions. In rows 1 to 3 (right), we show samples where the central feature vector, shown by the red marker, lies on a background region. In these cases, we still see sharp contours around the foreground objects, or around the body of water in row 1 (right). In cases like rows 4 to 8 (right), when there are multiple overlapping instances of the same object class, we see a uniform highlighting of the multiple object instances. We also see that DINO+LiFT better highlights thin structures in objects, like the teapot handle and tripod legs in rows 7 and 8 (left). For comparison, when DINO (without LiFT) is given a doubled input size, these details are sometimes lost to noisy background regions.
Table 12: Comparison between LiFT and other baselines on the Keypoint Correspondence task on SPair-71k for all metrics.
PCK@0.1 PCK@0.05 PCK@0.01
Method/Resolution 56 112 224 448 56 112 224 448 56 112 224 448
DINO 2.04 12.67 24.76 28.6 0.51 3.61 9.54 15.33 0.01 0.2 0.54 1.4
Leopart 2.35 11.2 23.33 26.54 0.6 3.22 8.9 12.26 0.05 0.1 0.47 0.79
DINO+LiFT 5.05 17.72 28.68 31.38 1.19 6.29 14.72 18.90 0.06 0.29 0.91 2.52
@@ Appendix 0.E Additional Similarity Map Samples
Method/Resolution 56 112 224 448 56 112 224 448 56 112 224 448
DINO 0.10 0.22 0.38 0.52 0.05 0.13 0.28 0.50 0.07 0.18 0.33 0.51
DINO+Bilinear 0.14 0.29 0.43 0.54 0.08 0.18 0.31 0.52 0.11 0.24 0.37 0.53
Leopart 0.09 0.20 0.35 0.47 0.05 0.12 0.26 0.43 0.07 0.16 0.30 0.45
SelfPatch 0.10 0.22 0.38 0.53 0.05 0.13 0.28 0.50 0.07 0.17 0.33 0.51
DINO+LiFT 0.16 0.33 0.48 0.59 0.10 0.23 0.41 0.63 0.13 0.28 0.44 0.61

### v2

@@ 6 Properties of LiFT
LiFT and Scale Invariance of Features. In this experiment, we demonstrate that LiFT intrinsically learns to generate features that are more scale-invariant. We use the Centered Kernel Alignment (CKA) metric [14, 32, 49], which can measure the similarity of a pair of feature maps even when they are different sizes. Using images in SPair-71k, we re-scale each image to a range of different sizes and then extract features with DINO or DINO+LiFT and measure the CKA similarity between all input size pairings. As a baseline, we also compare with bilinearly upsampled DINO features. In Figure 4, we take four source scales and plot the CKA similarity with the features at all other scales. We see that LiFT greatly improves the inter-scale feature similarity for small input scales, and moderately improves the similarity for medium and larger scales. This shows that LiFT produces representations that are more scale-invariant than the base DINO features. Bilinear upsampling does provide a small improvement in CKA similarity across scales, though the benefit is far smaller. This scale invariance property of LiFT is learned automatically, and likely comes from its multi-scale reconstruction objective. LiFT must learn to counteract the effect of input scale to map features from low resolution inputs to those of high resolution inputs. This property is desirable for dense tasks where objects appear at different scales.
Enhanced Self-Similarity Maps with LiFT. To further improve our understanding of the dense features generated by LiFT, we visualize the self-similarity of the features in Figure 5. Base DINO S/16 yields $14\times 14$ features for a $224\times 224$ image. In our comparisons, we visualize these features alongside: a bilinearly interpolated upsampling of these same features to $28\times 28$ (DINO+Bilinear), DINO features for a $448\times 448$ input image, and finally DINO+LiFT features generated for a $224\times 224$ input image. The last three configurations all yield $28\times 28$ feature grids. To visualize these features, we select the center-most token feature for each of the maps and compute its similarity with all other features and visualize the similarity scores. We show six feature similarity maps in Figure 5, and additional sample visualizations can be found in Appendix E. We find that, qualitatively, the LiFT output gives a cleaner boundary and better highlights the content corresponding to the central patch. Having a high similarity to relevant regions and a clear boundary is beneficial for multiple localized downstream tasks. It also indicates that the DINO+LiFT features have better spatial awareness of object boundaries. Note that for row 2, where the central pixel corresponds to the background region, the similarity map highlights the background in the image, though it still appears that DINO+LiFT produces better similarity maps with sharper edges. In row 3, when there are multiple objects of the same type, DINO+LiFT better highlights the separate object instances. These results are qualitative, but they suggest that the LiFT-enhanced features have better content and boundary information than the base DINO features, which likely contributes to their improved performance in correspondence and segmentation tasks.
@@
Variations in Backbone. One of the requirements of a feature densifying approach is that it should be easy to train on any backbone. To this end, we show that our approach consistently gives a performance gain with multiple different backbones: DINO (Table 8 row 1 vs. row 2), MoCo (Table 8 row 4 vs. row 5) and a fully-supervised ViT (Table 8 row 7 vs. row 8). This consistent improvement shows that LiFT can be trained in the exact same manner on differently trained ViTs without need for careful hyperparameter tuning. To verify that LiFT does not simply learn a bilinear interpolation, we apply LiFT modules trained on one backbone to the output of a different backbone. We show these results in Table 7 using the PCK@0.1 metric and $448\times 448$ images. It can be seen that when LiFT is applied to a different model than what was used to train it, the performance drops and is lower than not applying LiFT. This shows that LiFT learns a model-specific feature-densifying transform and not a simple interpolation.
@@
Figure 6: With no extra training, repeated application of LiFT produces pixel-dense feature maps.
Repeated Application of the LiFT Module. In our base approach, the LiFT module is applied once to the ViT features to double their resolution. In this section, we also test the potential benefits of applying the LiFT block multiple times. To check this, we take the LiFT module and super-resolute the features twice by passing the output of first super-resolution again through the same LiFT network. We denote this approach as ‘2$\times$×LiFT’ in Table 8. We can see that in most cases the performance increases after recursively applying LiFT. This is especially true for the $56\times 56$, $112\times 112$, and $224\times 224$ resolutions. For $448\times 448$, there are a few places where it negligibly drops performance, but it still shows improvement for most cases. Repeated application of LiFT can also be used to create detailed, pixel-dense feature self-similarity maps, as shown in Figure 8 for a 4×4$\times$4 × repeated application of the same pretrained LiFT module. Such maps could be useful for tasks like Unsupervised Semantic Segmentation [24].
@@ 7 Conclusion and Discussion
We have presented LiFT, a simple yet effective self-supervised Lightweight Feature Transform to boost the density of features of pretrained ViT backbones. This approach allows us to extract higher resolution spatial features from ViTs which can then be used for multiple dense downstream tasks. LiFT is task-agnostic and gives significant boosts in SPair Keypoint Correspondence, DAVIS Video Object Segmentation, Unsupervised Object Discovery, and COCO Detection and Segmentation. This benefit comes for a fraction of the computational cost compared with other densification methods. LiFT is a lightweight module that is easily trained with a self-supervised objective, and it is far cheaper than full backbone finetuning, as is done by other prior works. Through extensive experiments, we have shown that LiFT can be trained easily on any backbone and consistently leads to improved performance by generating better quality, higher-density features. We also show that LiFT can be applied to its own output in a recursive manner enabling good performance with even lower image resolutions. Finally, we show that our surprisingly simple method has several desirable emergent properties including scale-invariant features, and better object boundary maps. This makes LiFT a useful multipurpose tool for many potential downstream applications.
Acknowledgements. This work was partially supported by NSF CAREER Award (#2238769) to AS, and the NSF and NIST Institute for Trustworthy AI in Law and Society (TRAILS) (#2229885). The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. The authors acknowledge UMD’s supercomputing resources made available for conducting this research. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of NSF, NIST, or the U.S. Government. We would also like to thank our colleagues Matthew Gwilliam and Pravin Nagar for their feedback on this work.
@@ 0.A.4 Ablation of Training Epochs
Appendix 0.B Additional Details for ViTDet+LiFT
For our experiments combining LiFT with ViTDet [33], we increase the size of our LiFT module to address the additional complexity of the task and backbone. To be consistent with ViTDet, we use an MAE-trained ViT-Base backbone instead of the ViT-Small used in our other primary experiments. Note that a standard ViT-Small model outputs feature maps with $384$ channels, while ViT-Base outputs $768$ channels. To handle the increased number of channels, we commensurately increase the number of channels in the layers of our LiFT module. We also add an additional convolutional block to the encoder segment. This larger LiFT module has a total of $7M$ parameters, as compared with the $1.2M$ parameter version used for smaller architectures. The ViTDet model used has $111M$ parameters, so our combined ViTDet+LiFT architecture has $118M$ parameters total. This is a $6.3\%$ increase in total parameters, which is similar to the relative percentage increase of the smaller LiFT version for DINO ViT-S/16. Also, here we train LiFT on the COCO dataset in place of ImageNet. Because the COCO dataset is much smaller than ImageNet, we train on it for 100 epochs.
@@ Appendix 0.C Additional Backbones with LiFT
We further demonstrate the general utility of LiFT by applying it to several additional backbones, including Leopart [64] and several other DINO [6] ViTs, namely ViT-S/8, ViT-B/16, and ViT-B/8. The results are summarized in Table 11. LiFT shows consistent improvement for the various architectures and models across patch sizes (8 and 16), trainings (Leopart and DINO), and backbone sizes (Base and Small). We also extend the Performance vs. Compute Cost analysis curve to include both the MOCO and fully-supervised ViT backbones, as shown in Figure 7. We find that LiFT consistently boosts the performance of all three backbones at all FLOP allowances.
@@
Dataset Method Resolution
56 112 224 448
VOC07 DINO 20.74 50.07 65.60 68.27
Leopart 18.92 32.59 51.59 48.79
SelfPatch 18.04 41.99 62.40 63.62
DINO+BL 21.27 46.96 64.70 68.37
DINO+RC 28.96 55.00 66.85 68.87
DINO+JBU 26.16 56.60 66.75 69.03
DINO+LiFT 36.54 62.02 68.79 69.65
[…]
Leopart 22.44 37.41 55.74 54.40
SelfPatch 20.19 47.32 68.02 66.48
DINO+BL 23.46 52.64 68.53 71.55
DINO+RC 31.63 59.96 68.87 71.47
DINO+JBU 28.97 61.67 69.13 71.54
DINO+LiFT 40.56 66.21 70.91 71.71
[…]
Leopart 16.14 26.78 43.89 44.08
SelfPatch 14.15 35.76 52.18 55.47
DINO+BL 17.78 35.62 51.53 56.84
DINO+RC 22.92 42.53 54.52 58.40
DINO+JBU 21.36 43.87 55.45 58.82
DINO+LiFT 27.72 50.20 58.03 60.50
Appendix 0.D Additional Results
We present results for additional metrics on DAVIS, reporting the J Mean and F Mean in Table 13. We present these results alongside the previously reported J & F mean for completeness. We find that both the J Mean and F Mean are also consistently improved by adding LiFT, and that DINO+LiFT surpasses all other baselines. We also present additional Unsupervised Object Discovery results for PASCAL VOC 2007 [20] and PASCAL VOC 2012 [21] in Table 12 alongside the results for COCO20K. We again find that LiFT gives the best CorLoc boost over all baselines for both datasets.
@@ Appendix 0.D Additional Results
Method/Res. 56 112 224 448 56 112 224 448 56 112 224 448
DINO 9.50 21.90 37.80 52.10 5.20 13.10 28.10 49.70 7.40 17.50 33.00 50.90
Leopart 9.00 20.20 34.90 47.30 4.80 12.10 25.70 42.80 6.92 16.12 30.33 45.08
SelfPatch 9.70 21.60 38.10 52.50 5.10 12.90 27.90 50.30 7.40 17.23 33.01 51.37
[…]
DINO+RC 13.80 29.60 43.00 53.90 8.20 18.40 31.90 52.50 11.00 24.00 37.40 53.20
DINO+JBU 13.90 30.60 42.80 54.90 8.60 21.70 35.10 54.10 11.20 26.20 39.00 54.50
[…]
Appendix 0.E Additional Similarity Map Samples
We have found that the feature self-similarity maps for DINO+LiFT more clearly and sharply outline the central object in an image. To further highlight this, we provide a zoomed-in comparison of the difference between DINO+LiFT and DINO+Bilinear upscaling in Figure 8. We can see that DINO with Bilinear upsampling highlights the main object, but the outline is hazier and less precise due to the smoothing of the features. Meanwhile, the upscaled feature map produced by LiFT better respects object contours and produces a much sharper feature self-similarity map. Finally, we provide additional samples further showing the benefits of LiFT for self-similarity maps, as shown in Figure 9. In rows 1 to 8 (left), we show samples with single central objects of differing shapes and sizes. We see that the feature self-similarity maps for DINO+LiFT more uniformly fill the foreground object region, and have less noisy correlations with background regions. In rows 1 to 3 (right), we show samples where the central feature vector, shown by the red marker, lies on a background region. In these cases, we still see sharp contours around the foreground objects, or around the body of water in row 1 (right). In cases like rows 4 to 8 (right), when there are multiple overlapping instances of the same object class, we see a uniform highlighting of the multiple object instances. We also see that DINO+LiFT better highlights thin structures in objects, like the teapot handle and tripod legs in rows 7 and 8 (left). For comparison, when DINO (without LiFT) is given a doubled input size, these details are sometimes lost to noisy background regions.